OpenStudy (anonymous):

@e.mccormick Use a matrix to solve the linear system of equations: {5x-2y=27, -3x+5y=18}

OpenStudy (e.mccormick):

K. Know how to put them into matrix form?

OpenStudy (anonymous):

No. I was told to tag you =) I don't know anything about matrices. My lesson doesn't explain it very well. I was hoping you could sort of teach it to me? =/ Hope I'm not asking too much.

OpenStudy (e.mccormick):

The matrix form for this uses what is called an augmented matrix. Basically it uses a | in place of the = signs and gets rid of the variables: \(\left[\begin{array}{cc|c} 5 & -2 & 27\\ -3 & 5 & 18 \end{array}\right]\)

OpenStudy (anonymous):

Oh, okay.

OpenStudy (e.mccormick):

Now, the thing about solving or reducing matrices is what they call elementary row operations. You can multiply a row by a constant. You can multiply a row by a constant and add it to another row. That is basically it.

OpenStudy (e.mccormick):

Once you get things to the form of: \(\left[\begin{array}{cc|c} 1 & 0 & x\\ 0 & 1 & y \end{array}\right]\) Whatever is in the x and y would be the answer.

OpenStudy (anonymous):

Okay. So how do I go about reducing it to that form?

OpenStudy (e.mccormick):

I'm doing up an example for the first operation.

OpenStudy (e.mccormick):

Let me take the first row: 5 -2 | 27 I need to multiply it by something to add it to the -3 5 | 18 row in such a way that it would get rid of the 3. Well, to turn 5 into 3 you multiply by \(\dfrac{3}{5}\). That multiplies through the entire row: \(5\cdot\dfrac{3}{5} -2\cdot\dfrac{3}{5} | 27\cdot\dfrac{3}{5}\implies \) \(3 -\dfrac{6}{5} |\dfrac{81}{5} \) So I am going to add that and the last row to replace the last row. Does that make sense?

OpenStudy (anonymous):

Ahh, so you have to use multiplication or addition to completely get rid of one of the rows?

OpenStudy (e.mccormick):

Yah, zero it.

OpenStudy (anonymous):

Okay. Give me a second to think on that =) I need to write it down and actually go through it myself to see how it works.

OpenStudy (e.mccormick):

\(\begin{array}{cc|c} 3 & -\dfrac{6}{5} & \dfrac{81}{5}\\ -3 & \dfrac{25}{5} & \dfrac{90}{5} \\ \hline 0 & \dfrac{18}{5} & \dfrac{171}{5} \end{array}\) Now, that would be my new row 2, but it is ugly.

OpenStudy (e.mccormick):

So I would multiply it by \(\dfrac{5}{18}\) before I put it in place.

OpenStudy (e.mccormick):

That got me to: \(\left[\begin{array}{cc|c} 5 & -2 & 27\\ 0 & 1 & \dfrac{19}{2} \end{array}\right]\) Now, it should be super easy to see how to eliminate the -2 in row 1 with that. \(\ddot \smile\) And that is basically how you solve systems in matrix form.

OpenStudy (anonymous):

Okay, I'm back. Give me a second to read all this again, please?

OpenStudy (anonymous):

Alright. I have a question. Why is it that in the second row, you multiply everything by 3/5 except 3? Shouldn't you do the same thing to everything?

OpenStudy (e.mccormick):

In the second row I multiplied two things by 1. My version of 1 eas 1 for the first one but \(\frac{5}{5}\) for the second and third.

OpenStudy (anonymous):

I see.

OpenStudy (e.mccormick):

1 has many faces. \(\frac{e}{e}\) is one. 0 also hides a lot. \(1-1\) is 0. So you can add 0 to anything and multiply anything by 1 and not change it. The adding 0 is useful for completing the square.

OpenStudy (anonymous):

Hmmm. This just keeps getting more complicated =)

OpenStudy (anonymous):

So I'm trying to take what you told me from the first problem and apply it to the second one I have to do.

OpenStudy (e.mccormick):

Well, this one still needs a little work.

OpenStudy (anonymous):

Oh, okay. *facepalm* I thought you were done, haha.

OpenStudy (e.mccormick):

\(\left[\begin{array}{cc|c} 5 & -2 & 27\\ 0 & 1 & \dfrac{19}{2} \end{array}\right]\) vs \(\left[\begin{array}{cc|c} 1 & 0 & x\\ 0 & 1 & y \end{array}\right]\) So you have the y, but not the x.

OpenStudy (anonymous):

Okay. So I'm assuming the top row on the left needs to match the top row on the right? How do I make that happen?

OpenStudy (e.mccormick):

Well, first you need to eliminate the -2. So what multiple of row 2 could you add to row 1 to do that?

OpenStudy (e.mccormick):

Seems you got bumped. Here are some notes on elementary row operations: http://www.purplemath.com/modules/mtrxrows.htm